the excess(equal in no) of electrons that must be placed on each of two spheres spaced 3 cm apart, with force of repulsion between sphere to be 10*-19 is

Grade:12

1 Answers

Arun
5 years ago
 

Dear pallavi

let the charge on spheres is q

so F= 1/4∏ε qq/.032

  given F= 10-19

       1/4∏ε qq/.03=10-19

 9 *109 * q2 /.03=10-19 

               q=10-16

1 electron has  1.6 * 10-19 C  

 ne =q

 n=q/e =10-16/1.6 * 10-19 C

           =625

 

Regards

Arun